OpenStudy (anonymous):

1 2/3 divided by 1 2/5

OpenStudy (lgbasallote):

5/3/7/5 = 25/21

OpenStudy (saifoo.khan):

First convert to improper fraction: \[\frac{3(1) +2}{3} \div \frac{5(1) +2}{5}\]\[\frac53 \div \frac75\]Flip action, \[\frac53 \times \frac57\]\

OpenStudy (saifoo.khan):

\[\frac{5 \times 5}{3 \times 7}\to \frac{25}{21}\]